Please help me solve this equation:
Given the sum of a and a3 = 3
and the sum of a4 and a6 = 81.
1) find the common ratio, r
p/s: This is what I've done. But then, I'm stuck.
= a+ar^3-1=3
= a+ar^2=3
= a(1+r^2)=3 << FACTORIZE.
= ar^4-1+ar^6-1=81
= ar^3+ar^5=81
= ar^3 (1+r^2)=81 << FACTORIZE.
then, what should I do?

Answer by lynnlo(4176)

A+AR^3-1=3
MOVE THE -3 TO THE LEFT
A+AR^3-1-3=0
ADD FRACTIONS OR YOUR INTEGERS TOGETHER
A+AR^3+-4=0
REMOVE THE PLUS SIGN(+) IN FRONT OF THE -4
A+AR^3-4=0
R=3√4/A-1